OpenStudy (anonymous):

can someone tell me how to prove for infinite limit with epsilon and delta or N and M?

OpenStudy (abb0t):

Yes, if you gave us an example, we could.

OpenStudy (abb0t):

What's the function?

OpenStudy (anonymous):

l\[\lim_{x \rightarrow \infty}3x^{2}-x-2/5x^{2}+4x+1\]

ganeshie8 (ganeshie8):

\[\lim_{x \rightarrow \infty}~\dfrac{3x^{2}-x-2}{5x^{2}+4x+1}\] like that ?

OpenStudy (anonymous):

Yes, like that

ganeshie8 (ganeshie8):

so what does it evaluate to ?

ganeshie8 (ganeshie8):

you need to know the limit before starting to prove it using epsilon-delta

OpenStudy (anonymous):

i know the limit it's 3/5

OpenStudy (anonymous):

but the prove is quite confusing.

ganeshie8 (ganeshie8):

its easy once you know how to play the game

OpenStudy (anonymous):

i guess... so :)

ganeshie8 (ganeshie8):

Let \(f(x) =\dfrac{3x^{2}-x-2}{5x^{2}+4x+1} \) given any positive number \(\epsilon\), you have to show that there exists some \(x\) value, after which the difference between \(f(x)\) and \(\dfrac{3}{5}\) is less than \(\epsilon\).

ganeshie8 (ganeshie8):

Given any positive number \(\epsilon\), prove that there exists some \(x\) value beyond which below is true : \(\left|\dfrac{3x^2-x-2}{5x^2+4x+1}-\dfrac{3}{5}\right|\lt \epsilon\)

ganeshie8 (ganeshie8):

simplify and try expressing \(x\) in terms of \(\epsilon\)

ganeshie8 (ganeshie8):

@kal1921

OpenStudy (anonymous):

simplify? sorry where is delta?

OpenStudy (zzr0ck3r):

Simplify so that you can find the bound on delta. You have to work backwards

ganeshie8 (ganeshie8):

yeah what we're doing is just scratch work to express \(x\) in terms of \(\epsilon\)

OpenStudy (anonymous):

\[-\epsilon< equation<\epsilon\] i get this...what is next?

ganeshie8 (ganeshie8):

for \(x\gt 1\), we have : \( \left|\dfrac{3x^2-x-2}{5x^2+4x+1}-\dfrac{3}{5}\right|\\~\\ =\left|\dfrac{17x+13}{5(5x^2+4x+1)}\right|\\~\\ \lt\left|\dfrac{17x+13}{25x^2}\right|\\~\\ \lt\left|\dfrac{17x+17x}{25x^2}\right|\\~\\ =\left|\dfrac{34}{25x}\right|\\~\\ \) \(\dfrac{34}{25x} \lt \epsilon \implies x \gt \dfrac{34}{25\epsilon}\) therefore \(x = \dfrac{34}{25\epsilon}\) works, now you can start the proof.

OpenStudy (anonymous):

sorry, how can i relate x and delta... if it x wasn't approaching to infinity i would say /x-a/>delta

ganeshie8 (ganeshie8):

The definition for limits at infinity is slightly different : \(\lim\limits_{x\to\infty}f(x)=L\) means that for "every" \(\epsilon\gt 0\), there exists some \(x\) value, \(N\), such that \(|f(x)-L|\lt\epsilon\) for all \(x\gt N\).

ganeshie8 (ganeshie8):

In present problem, we have cooked up that \(N\) to be \(\dfrac{34}{25\epsilon}\)

ganeshie8 (ganeshie8):

you choose "any" positive \(\epsilon\), the value \(|f(x)-\frac{3}{5}|\) will always be less than that chosen \(\epsilon\) whenever \(x\) is greater than \(\dfrac{34}{25\epsilon}\)

ganeshie8 (ganeshie8):

In other words, we can make \(f(x)\) as close to \(\frac{3}{5}\) as we want by making \(x\) sufficiently large.

ganeshie8 (ganeshie8):

I think this is same as the N and M definition that you were refering to in main question